OpenStudy (anonymous):

Which of the following is the simplified form of 4y^2 - 1 / 2y^2 + 3y - 2?

OpenStudy (anonymous):

is this / as in divide or u use it as a comma?

OpenStudy (anonymous):

means divide

OpenStudy (anonymous):

try to factorize numerator and denominator

OpenStudy (paxpolaris):

\[\Large{4y^2-1 \over 2y^2+3y-2}\] Now, \[4y^2-1=\left( 2y+1 \right)\left( 2y-1 \right)\]\[2y^2+3y-2= (2y-1)(y+2)\]

OpenStudy (paxpolaris):

\[\Large {\left( 2y+1 \right) \cancel {\left( 2y-1 \right)}\over \cancel {\left( 2y-1 \right)}\left( y+2 \right) }\]